We are required to form different words with the help of letters of the word 'INTEGER'.
let m1 be the number of words in which 'I' and 'N' are never together and m2 be the number of words which begin with 'I 'and end with 'R',
then m1/m2 is given by ?

Answer by khwang(438)

All possible words: 7!/2! = 7*6!/2 (since there are 2 E's)
If TN together TN or NT as a letter, the other five are EE,I,G,R
so there are 2*6!/2! = 6! = 720
Hence,m1 = 7*6!/2 - 720 = 720(7/2 -1) = 720 (5/2) = 1800

m2 = 5!/2 = 60 (why?)
So, m1/m2 = 30

Kenny